Complete the square to determine whether the equation represents an ellipse, a parabola. If the graph is an ellipse, find the center, foci, vertices, and lengths of the major and minor axes. If it is a parabola, find the vertex, focus, and directrix. Then sketch the graph of the equation.

4x^2 +4x − 8y + 9 = 0
Answer #1

Ч Y +yu -8y + g =0 x + x) = 8y-g = 4(**+n++) = 8y-19+1 = 8y - 8 » ul^+ )? 814-1) (n + 1)² = 2(y!) = 4 (2) lyri) fence, Given